OpenStudy (anonymous):

Hi, do you know how to solve for y. The eqution is x=(2y-1)/(y-3)

OpenStudy (anonymous):

x(y-3) = 2y-1 xy - 2y = 3x - 1 y(x-2)=3x-1 y=(3x-1)/(x-2)